# Relativistic structure of a supermassive black hole embedded in the dark matter halo of NGC 4649 (M60)

### v1

@@ IV Conclusion
From a geometrical perspective, the solution features a single event horizon, the size of which is modulated by the DM halo via the influence of the three aforementioned parameters. We examined the Kretschmann scalar $K$, a curvature invariant used to diagnose the presence and nature of spacetime singularities. The analysis reveals a divergent behavior as $r\to 0$, signaling the presence of a central singularity, and an asymptotic flatness as $r\to\infty$. In the intermediate regime, the curvature structure is significantly altered by the halo, indicating that the mass distribution induced by the halo strongly shapes the geometry of the surrounding spacetime.
A key observational feature of BHs, namely their shadow, was also examined in this context. Building on methodologies established in [37], we analyzed how the radius of the BH shadow, $r_{sh}$, responds to variations in halo parameters. Our results show that $r_{sh}$ is particularly sensitive to changes in the halo scale radius $a$, while the critical velocity $V_{c}$ affects the shadow indirectly by influencing $a$. These findings are consistent with prior modeling in [36], and point to the shadow as a potentially powerful observational window into the DM environments of astrophysical BHs.
In the thermodynamic domain, our analysis focused on how the presence of the DM halo modifies traditional relations, particularly the mass-entropy relation $M(S)$. Despite the additional mass contributions from the halo, the relation $M^{2}\propto S$ is approximately preserved due to the low values of $V_{c}$ adopted in our model. Notably, the mass function shows a decreasing trend relative to the Schwarzschild case, such that for high entropy values, the BH mass asymptotically approaches a lower bound: $M\sim(1-2V_{c}^{2})M^{(\text{sch})}$ as $S\to\infty$. This behavior illustrates how the halo effectively alters the energy budget of the BH.

### v2

@@ III.3 Black Hole Shadow Radius
In Tables 3 and 4, we apply the same approach as in expressions (37) and (38) to highlight the subtle modifications to the BH shadow induced by variations in $V_{c}$ and $a$ within the parameter ranges specified in Table 1. Specifically, Table 3 examines the ratio $r_{sh}/m_{BH}$ under changes in $V_{c}$, while Table 4 explores the same ratio under variations in $a$.
Table 3 corresponds to the graphs shown in Fig. 5, demonstrating that although the shadow radius $r_{sh}$ varies with $V_{c}$, the values of ${\rm RatioI}(V_{c})$ and ${\rm RatioII}(V_{c})$ remain essentially constant. This indicates that within this interval, a non-zero $V_{c}$ is sufficient to produce an increase in $r_{sh}$. We note an apparent oscillation of ${\rm RatioI}(V_{c})$ and ${\rm RatioII}(V_{c})$ about stable averages $\overline{\rm RatioI}(V_{c})=0.932867\times 10^{-7}$ and $\overline{\rm RatioII}(V_{c})=1.08344\times 10^{-7}$, with corresponding standard deviations $\omega_{I}=0.0187499\times 10^{-7}$ and $\omega_{II}=0.0115171\times 10^{-7}$, respectively. Both standard deviations are small relative to their means, indicating low variability of the values reported in Table 3. Consequently, the coefficient of variation, given by $\frac{\omega_{I}}{\overline{\rm RatioI}(V_{c})}\times 100\%\approx 2\%$ and $\frac{\omega_{II}}{\overline{\rm RatioII}(V_{c})}\times 100\%\approx 1\%$, shows that ${\rm RatioII}(V_{c})$ is more stable than ${\rm RatioI}(V_{c})$. This observation aligns with [36], which suggests that the parameter values in Data II are more precise than those in Data I.
Table 4 corresponds to the graphs in Fig. 6, illustrating that both ratios increase with $a$, which suggests that an expansion of the halo’s extent leads to an increase in the shadow radius. At every value of $a$, the inequality ${\rm RatioI}(a)>{\rm RatioII}(a)$ holds, indicating that the Data I configuration exhibits a greater variation in $r_{sh}$ compared to Data II. This increased sensitivity to changes in the halo parameters aligns with the fact that $(V_{cII},a_{II})>(V_{cI},a_{I})$, reflecting a stronger response of spacetime curvature. These findings are consistent with the behaviors observed in Figs. 3 and 4.
$V_{c}$ ${\rm RatioI}(V_{c})[\times 10^{-7}]$ ${\rm RatioII}(V_{c})[\times 10^{-7}]$
$0.001$ $0.94926978944$ $1.0848654197$
$0.0015$ $0.91375943034$ $1.0859227717$
$0.002$ $0.90629899585$ $1.0676928408$
$0.0025$ $0.94801603900$ $1.0958764474$
$0.003$ $0.93327437556$ $1.0719221444$
$0.0035$ $0.94658315408$ $1.0943817941$
Table 3: We define the functions ${\rm RatioI}(V_{c})=\frac{r_{sh}(m_{BHI},V_{c},a_{I})/m_{BHI}}{r_{sh}^{(sch)}/M_{0}}-1$ and ${\rm RatioII}(V_{c})=\frac{r_{sh}(m_{BHII},V_{c},a_{II})/m_{BHII}}{r_{sh}^{(sch)}/M_{0}}-1$ using $r_{sh}$ from Eq. (36), and $r_{sh}^{(sch)}/M_{0}=3\sqrt{3}$ from the Schwarzschild solution without halo. The values of $V_{c}$ are consistent with Table 1.
$a$[m] ${\rm RatioI}(a)[\times 10^{-7}]$ ${\rm RatioII}(a)[\times 10^{-7}]$
$5\times 10^{20}$ $1.4972686757$ $1.1580407233$
$1\times 10^{21}$ $3.1029277814$ $2.4392042386$
$1.5\times 10^{21}$ $4.7085869159$ $3.6875174048$
$2\times 10^{21}$ $6.2297376657$ $4.7387284807$
$2.5\times 10^{21}$ $7.7508884422$ $6.3155451291$
Table 4: We define the functions ${\rm RatioI}(a)=\frac{r_{sh}(m_{BHI},V_{cI},a)/m_{BHI}}{r_{sh}^{(sch)}/M_{0}}-1$ and ${\rm RatioII}(a)=\frac{r_{sh}(m_{BHII},V_{cII},a)/m_{BHII}}{r_{sh}^{(sch)}/M_{0}}-1$ using $r_{sh}$ from Eq. (36), and $r_{sh}^{(sch)}/M_{0}=3\sqrt{3}$ from the Schwarzschild solution without halo. The values of $a$ are consistent with Tabela 1.
Equations (37) and (38), together with Tables 3 and 4, demonstrate that the difference in the shadow radius $r_{sh}$ between the halo solution and the Schwarzschild solution (without halo) is exceedingly small. Such a difference is experimentally challenging to detect given the precision limits of current observational instruments. In [53], the authors studied the shadow of the Sgr A* black hole under DM density profiles different from the one considered here, specifically for the Cold Dark Matter (CDM) and Scalar Field Dark Matter (SFDM) models. Their results indicate that the presence of the halo causes a slight increase in the shadow radius, approximately $9.8\times 10^{-3}\%$ for CDM and $9.8\times 10^{-5}\%$ for SFDM. However, these effects are only significant for extreme values of the parameter $k\sim\rho_{c}R^{3}\sim 10^{7}$, where $\rho_{c}$ is the central density and $R$ is the radius at which pressure and density vanish. In the present work, the calculated increases in $r_{sh}$ amount to approximately $0.926837\times 10^{-5}\%$ for Data I and $1.05949\times 10^{-5}\%$ for Data II. Therefore, although the presence of the dark matter halo does induce changes in the shadow radius, these variations are sufficiently subtle to be currently unobservable with existing measurement techniques.
@@ IV Conclusion
From a geometrical perspective, the solution features a single event horizon, the size of which is modulated by the DM halo via the influence of the three aforementioned parameters. We examined the Kretschmann scalar $K$, a curvature invariant used to diagnose the presence and nature of spacetime singularities. The analysis reveals a divergent behaviour as $r\to 0$, signalling the presence of a central singularity, and an asymptotic flatness as $r\to\infty$. In the intermediate regime, the curvature structure is significantly altered by the halo, indicating that the mass distribution induced by the halo strongly shapes the geometry of the surrounding spacetime. We evaluated the percentage variation in curvature as a function of radius using the ratio $K_{0}(r)/K(r)$, where we observed that in regions close to the center of the BH, curvature is influenced by the change in mass caused by the presence of the halo.
A key observational feature of BHs, namely their shadow, was also examined in this context. Building on methodologies established in [37], we analysed how the radius of the BH shadow, $r_{sh}$, responds to variations in halo parameters. Our results show that $r_{sh}$ is particularly sensitive to changes in the halo scale radius $a$ and the critical velocity $V_{c}$, both of which affect the shadow. These findings are consistent with prior modelling in [36], and with the conclusions obtained by [53]. Therefore, although the radius of the shadow is barely affected by the presence of the halo, measuring it can provide valuable information about the DM environment surrounding astrophysical BHs.
In the thermodynamic domain, our analysis focused on how the presence of the DM halo modifies traditional relations, particularly the mass-entropy relation $M(S)$. Despite the additional mass contributions from the halo, the relation $M^{2}\propto S$ is approximately preserved due to the low values of $V_{c}$ adopted in our model. Notably, the mass function shows a decreasing trend relative to the Schwarzschild case, such that for high entropy values, the BH mass asymptotically approaches a lower bound: $M\sim(1-2V_{c}^{2})M^{(\text{sch})}$ as $S\to\infty$. This behaviour illustrates how the halo effectively alters the energy budget of the BH.